johnny walked from home to school at rate of 3mph. he came back home at rate of 5mph. total trip took him 1 hour. how many miles did he walk?
Answer by Alan3354(69443)

johnny walked from home to school at rate of 3mph. he came back home at rate of 5mph. total trip took him 1 hour. how many miles did he walk?
-----------
The avg speed of a round-trip is
2*3*5/(3+5) = 30/8 mi/hr
= 3.75 mi/hr
3.75*1 = 3.75 miles round trip
-----------------
I used the avg speed calculation that I worked out some time ago.
You can accept it, or not.